# EDVD-LLaMA: Explainable Deepfake Video Detection via Multimodal Large Language Model Reasoning

### v1

@@ III-E Construction Process of ER-FF++set
Motivation: Existing deepfake video datasets only provide binary classification labels for the video modality, lacking visual-language samples suitable for training MLLM in the DVD task. This limitation makes it difficult for MLLM to understand fine-grained instructions related to deepfake details, thereby hindering the development of interpretable reasoning frameworks. Constructing an interpretable reasoning dataset requires leveraging pretrained MLLM to transform the manipulation information in the visual modality of existing DVD datasets into precise textual descriptions. The key challenge in this transformation process lies in suppressing the hallucinated outputs of pretrained MLLM to prevent noise from contaminating the constructed dataset. To address this issue, our core contributions are focused on two aspects: (1) For each deepfake technique, we designed targeted prompts based on its unique characteristics to guide MLLM (like Qwen2.5-VL [42]) in focusing on different deepfake features and providing more accurate visual cues. (2) We provided MLLM with consecutive frames of deepfake videos along with their corresponding masks and facial landmarks, enabling MLLM to generate rationales that focus on manipulated regions, thereby suppressing hallucinated outputs and delivering more accurate deepfake cues.
@@
Data Collection: Based on [18], we categorize common video forgeries into five types: DeepFake [43], Face2Face [44], FaceSwap [45], FaceShifter [46] and NeuralTexture [47]. As shown in step 1 of Fig. 4, we collect forged videos of these five types, along with their corresponding real videos and self-constructed data, from public datasets.
Assistant MLLM-Aided Rationale Generation: Manual analysis of forged videos is time-consuming and inefficient. Therefore, we employ MLLM[42] as the assistant LLM for automated analysis of forged videos. The output analysis follows a structured format including detection results, reasoning process, and localization description.
@@
In step 3 of Fig. 4, we first generate forgery region masks [48] by pixel-level comparison between real and forged video frames. Subsequently, the video is divided into regions such as the mouth, nose, eyes, and face based on facial landmarks. The degree of forgery in each area is quantitatively assessed to form the structured labels. These structured labels are highly accurate and traceable, effectively reducing hallucinations that may occur during automatic text generation by MLLM. We input the structured labels, region localization information, and dedicated prompts for different forgery techniques into the assistant LLM. For different types of forgery, the LLM conducts targeted analysis based on the structured annotations, quantitatively assesses the degree of forgery in each region, and provides standardized annotations of the tampering method along with rationales that include structured localization information.
The ER-FF++ set contains a total of 7,362 video clips, which are divided according to the proportions shown in the Overview section of Fig. 4. In terms of authenticity distribution, there are 5,000 Deepfake videos (67.9%) and 2,362 Real videos (32.1%). The forgery types cover five mainstream methods (DeepFake [43], Face2Face [44], FaceSwap [45], FaceShifter [46] and NeuralTexture [47]), with each type being evenly distributed. This balanced setting of forgery types and reasonable dataset partitioning help models learn generalizable and interpretable reasoning and detection capabilities under various manipulation mechanisms, and support the evaluation of cross-method generalization and fine-grained reasoning abilities.
@@ IV-A Experiment Setup
Dataset: We construct the training and test sets of ER-FF++set using the dataset construction method. For the training set, we use videos manipulated by Deepfake [43], Face2Face [44], FaceSwap [45], FakeShifter [46], and NeuralTexture [47] techniques, as well as real videos subjected to data augmentation, as the source data. For the test set, in addition to the test portion of ER-FF++set, we also select several challenging public benchmark datasets, including Celeb-DF [7] and WildDeepfake [49].
@@
State-of-the-Art Methods: To ensure a fair comparison, we selected competitive methods that provide open-source code or pretrained models. To evaluate the DVD performance of EDVD-LLaMA, we compared it with traditional methods such as Xception [50], DSP-FWA [7], B4 [51], B4ATT [52], MCX-API [53], and S-TAD [54]. Additionally, we compared it with open-source MLLMs, including Video-LLaVA [55], Video-ChatGPT [56], VideoChat [32], and VideoLLaMA3 [57]. All these methods were retrained on ER-FF++set to ensure consistency with the same training settings. Furthermore, to assess the explainability of EDVD-LLaMA, we compared it with the aforementioned open-source MLLMs (VideoLLaVA [55], Video-ChatGPT [56], VideoChat [32], and VideoLLaMA3 [57]).
@@
Evaluation Metrics: For detection, we report video frame-level accuracy (Acc.), AUC, and F1 score. To evaluate explainability, we adopt multiple automated evaluation metrics, including CIDEr [58], ROUGE_L [59], BLEU-4 [60], METEOR [61], and Cosine Semantic Similarity (CSS). These metrics comprehensively measure the similarity between generated and reference texts from various perspectives, including term frequency–inverse document frequency (TF-IDF) weighting, longest common subsequence, n-gram overlap, synonym matching, and semantic vector similarity. For EDVD-LLaMA, the default threshold of 0.5 is applied unless otherwise specified.
@@
Implementation Details: On ER-FF++set, we initially fine-tuned the $\text{LLM}^{(1)}$ (Qwen2.5-7B) [16] using LoRA (rank=128, alpha=256), while applying full-parameter training to certain components of the ST-SIT module, such as the feature fusion and bridging layers. The training was conducted for 4 epochs on four NVIDIA RTX 6000 Ada 48G GPUs, with a learning rate of 2e-5 and a batch size of 1. Subsequently, we fine-tuned the $\text{LLM}^{(2)}$ (Qwen2.5-7B) [16] with LoRA using the same configuration, training for 8 epochs on the same hardware setup, with a learning rate of 1e-5 and a batch size of 1. During training, we set the composite loss weights to $\lambda_{1}=0.8$, $\lambda_{2}=1$, and $\lambda_{3}=0.1$ to balance rationale learning, answer prediction, and evidence consistency.
IV-B Comparsion with MLLMS
@@
To evaluate the video deepfake detection capability of our proposed method, we compare EDVD-LLaMA with several representative MLLMs on the ER-FF++set, as shown in Tab. I. The results demonstrate that EDVD-LLaMA achieves the best performance across all metrics, with an Acc. of 84.7, AUC of 82.6, and F1 score of 81.1. These results are significantly higher than those of the other MLLMs, such as Video-LLaVA, Video-ChatGPT, and VideoChat, whose accuracy and F1 scores are close to random guessing. Compared to the advanced VideoLLaMA3, our model still surpasses it by a large margin in all metrics (nearly 12% average improvement). This highlights the superior effectiveness of EDVD-LLaMA for video deepfake detection, benefiting from its tailored feature extraction and interpretable reasoning framework.
@@
TABLE I: Performance comparison of EDVD-LLaMA with MLLMs on the ER-FF++set for video deepfake detection.
@@
To evaluate the quality of explanatory text generation, we compare EDVD-LLaMA with several representative MLLMs on the ER-FF++set using CIDEr, ROUGE_L, BLEU-4, METEOR, and CSS metrics. As shown in Tab. II, EDVD-LLaMA achieves the highest performance across all metrics. Specifically, EDVD-LLaMA obtains a CIDEr score of 2.4579, outperforming VideoLLaMA3 (2.3154), Video-ChatGPT (2.2416), VideoChat (2.1952), and Video-LLaVA (2.1738). Similarly, EDVD-LLaMA achieves the best ROUGE_L (0.7013), BLEU-4 (0.5048), METEOR (0.4236), and CSS (0.7135), consistently surpassing all baselines. These results demonstrate that EDVD-LLaMA generates more accurate, fluent, and semantically consistent explanatory texts for video deepfake detection compared to existing MLLMs.
@@ IV-C Generalization Performance Cross Forgery Methods and Datasets
To further evaluate the generalization ability of different methods across various forgery types, we conducted cross-forgery experiments. As shown in Tab. IV, we performed cross-testing among different forgery types in the ER-FF++ dataset, where models are trained on one type of forgery and tested on the others. The results demonstrate that traditional DVD methods exhibit limited performance when transferring across different types of forgery. In contrast, our proposed method consistently achieves the best average score across forgery methods, indicating superior cross-type generalization capability.
TABLE IV: Cross-forgery evaluation results on the ER-FF++ dataset. Models are trained on one type of forgery and tested on others. Cross Avg. denotes the average score across forgery methods.
Train
Methods Test(Intra Dataset——Cross Forgery)
Cross
Avg.
@@
Face2Face
Xception 56.28 - 50.73 49.74 51.43 52.05
@@ IV-D Qualitative results
In Fig. 5, we present a qualitative comparison of EDVD-LLaMA with other baseline models (such as Video-LLaVA [55], Video-ChatGPT [56], VideoChat [32], and VideoLLaMA3 [57]) on the ER-FF++ dataset. This dataset contains samples of deepfake videos and real videos, where the models are required to determine the authenticity of each video. EDVD-LLaMA demonstrates outstanding performance in detecting deepfake videos through its unique design. The results show that EDVD-LLaMA can accurately identify subtle artifacts in deepfake videos, such as decoupled motion vectors and anomalies in dynamic changes. In contrast, other baseline models often fail to handle these complex scenarios due to insufficient sensitivity to artifacts or incomplete feature extraction, resulting in misjudgments. For example, VideoLLaVA and Video-ChatGPT fail to detect critical artifacts in certain deepfake videos, leading to incorrect predictions, while VideoChat and VideoLLaMA3 occasionally generate false positives when processing real videos. In conclusion, EDVD-LLaMA not only achieves accurate judgments of video authenticity but also provides detailed reasoning, proving the effectiveness and potential of its design. More qualitative results of our model can be found in the supplementary video.

### v2

@@ III-E Construction Process of ER-FF++set
Motivation: Existing deepfake video datasets only provide binary classification labels for the video modality, lacking visual-language samples suitable for training MLLM in the DVD task. This limitation makes it difficult for MLLM to understand fine-grained instructions related to deepfake details, thereby hindering the development of interpretable reasoning frameworks. Constructing an interpretable reasoning dataset requires leveraging pretrained MLLM to transform the manipulation information in the visual modality of existing DVD datasets into precise textual descriptions. The key challenge in this transformation process lies in suppressing the hallucinated outputs of pretrained MLLM to prevent noise from contaminating the constructed dataset. To address this issue, our core contributions are focused on two aspects: (1) For each deepfake technique, we designed targeted prompts based on its unique characteristics to guide MLLM (like Qwen2.5-VL [Qwen2.5-VL]) in focusing on different deepfake features and providing more accurate visual cues. (2) We provided MLLM with consecutive frames of deepfake videos along with their corresponding masks and facial landmarks, enabling MLLM to generate rationales that focus on manipulated regions, thereby suppressing hallucinated outputs and delivering more accurate deepfake cues.
@@
Data Collection: Based on [roessler2019faceforensicspp], we categorize common video forgeries into five types: DeepFake [Deepfakes2021], Face2Face [thies2016face2face], FaceSwap [FaceSwap2021], FaceShifter [li2019faceshifter] and NeuralTexture [thies2019deferred]. As shown in step 1 of Fig. 4, we collect forged videos of these five types, along with their corresponding real videos and self-constructed data, from public datasets.
Assistant MLLM-Aided Rationale Generation: Manual analysis of forged videos is time-consuming and inefficient. Therefore, we employ MLLM[Qwen2.5-VL] as the assistant LLM for automated analysis of forged videos. The output analysis follows a structured format including detection results, reasoning process, and localization description.
@@
In step 3 of Fig. 4, we first generate forgery region masks [sun2025towards] by pixel-level comparison between real and forged video frames. Subsequently, the video is divided into regions such as the mouth, nose, eyes, and face based on facial landmarks. The degree of forgery in each area is quantitatively assessed to form the structured labels. These structured labels are highly accurate and traceable, effectively reducing hallucinations that may occur during automatic text generation by MLLM. We input the structured labels, region localization information, and dedicated prompts for different forgery techniques into the assistant LLM. For different types of forgery, the LLM conducts targeted analysis based on the structured annotations, quantitatively assesses the degree of forgery in each region, and provides standardized annotations of the tampering method along with rationales that include structured localization information.
The ER-FF++ set contains a total of 7,362 video clips, which are divided according to the proportions shown in the Overview section of Fig. 4. In terms of authenticity distribution, there are 5,000 Deepfake videos (67.9%) and 2,362 Real videos (32.1%). The forgery types cover five mainstream methods (DeepFake [Deepfakes2021], Face2Face [thies2016face2face], FaceSwap [FaceSwap2021], FaceShifter [li2019faceshifter] and NeuralTexture [thies2019deferred]), with each type being evenly distributed. This balanced setting of forgery types and reasonable dataset partitioning help models learn generalizable and interpretable reasoning and detection capabilities under various manipulation mechanisms, and support the evaluation of cross-method generalization and fine-grained reasoning abilities.
@@ IV-A Experiment Setup
Dataset: We construct the training and test sets of ER-FF++set using the dataset construction method. For the training set, we use videos manipulated by Deepfake [Deepfakes2021], Face2Face [thies2016face2face], FaceSwap [FaceSwap2021], FakeShifter [li2019faceshifter], and NeuralTexture [thies2019deferred] techniques, as well as real videos subjected to data augmentation, as the source data. For the test set, in addition to the test portion of ER-FF++set, we also select several challenging public benchmark datasets, including Celeb-DF [li2020celeb] and WildDeepfake [zi2020wilddeepfake].
@@
State-of-the-Art Methods: To ensure a fair comparison, we selected competitive methods that provide open-source code or pretrained models. To evaluate the DVD performance of EDVD-LLaMA, we compared it with traditional methods such as Xception [chollet2017xception], DSP-FWA [li2020celeb], B4 [tan2019efficientnet], B4ATT [bonettini2021video], MCX-API [Xu_2023_WACV], and S-TAD [gao2024texture]. Additionally, we compared it with open-source MLLMs, including Video-LLaVA [lin2023video], Video-ChatGPT [maaz2024videochatgptdetailedvideounderstanding], VideoChat [li2024videochatchatcentricvideounderstanding], and VideoLLaMA3 [damonlpsg2025videollama3]. All these methods were retrained on ER-FF++set to ensure consistency with the same training settings. Furthermore, to assess the explainability of EDVD-LLaMA, we compared it with the aforementioned open-source MLLMs (VideoLLaVA [lin2023video], Video-ChatGPT [maaz2024videochatgptdetailedvideounderstanding], VideoChat [li2024videochatchatcentricvideounderstanding], and VideoLLaMA3 [damonlpsg2025videollama3]).
@@
Evaluation Metrics: For detection, we report video frame-level accuracy (Acc.), AUC, and F1 score. To evaluate explainability, we adopt multiple automated evaluation metrics, including CIDEr [Vedantam2015CIDEr], ROUGE_L [Lin2004ROUGE], BLEU-4 [Papineni2002BLEU], METEOR [Denkowski2014METEOR], and Cosine Semantic Similarity (CSS). These metrics comprehensively measure the similarity between generated and reference texts from various perspectives, including term frequency–inverse document frequency (TF-IDF) weighting, longest common subsequence, n-gram overlap, synonym matching, and semantic vector similarity. For EDVD-LLaMA, the default threshold of 0.5 is applied unless otherwise specified.
@@
Implementation Details: On ER-FF++set, we initially fine-tuned the $\text{LLM}^{(1)}$ (Qwen2.5-7B) [qwen2025qwen25technicalreport] using LoRA (rank=128, alpha=256), while applying full-parameter training to certain components of the ST-SIT module, such as the feature fusion and bridging layers. The training was conducted for 4 epochs on four NVIDIA RTX 6000 Ada 48G GPUs, with a learning rate of 2e-5 and a batch size of 1. Subsequently, we fine-tuned the $\text{LLM}^{(2)}$ (Qwen2.5-7B) [qwen2025qwen25technicalreport] with LoRA using the same configuration, training for 8 epochs on the same hardware setup, with a learning rate of 1e-5 and a batch size of 1. During training, we set the composite loss weights to $\lambda_{1}=0.8$, $\lambda_{2}=1$, and $\lambda_{3}=0.1$ to balance rationale learning, answer prediction, and evidence consistency.
IV-B Comparison with MLLMs
@@
To evaluate the deepfake video detection capability of our proposed method, we compare EDVD-LLaMA with several representative MLLMs on the ER-FF++set, as shown in Tab. I. The results demonstrate that EDVD-LLaMA achieves the best performance across all metrics, with an Acc. of 84.7, AUC of 82.6, and F1 score of 81.1. These results are significantly higher than those of the other MLLMs, such as Video-LLaVA, Video-ChatGPT, and VideoChat, whose accuracy and F1 scores are close to random guessing. Compared to the advanced VideoLLaMA3, our model still surpasses it by a large margin in all metrics (nearly 12% average improvement). This highlights the superior effectiveness of EDVD-LLaMA for deepfake video detection, benefiting from its tailored feature extraction and interpretable reasoning framework.
@@
TABLE I: Performance comparison of EDVD-LLaMA with MLLMs on the ER-FF++set for deepfake video detection.
@@
To evaluate the quality of explanatory text generation, we compare EDVD-LLaMA with several representative MLLMs on the ER-FF++set using CIDEr, ROUGE_L, BLEU-4, METEOR, and CSS metrics. As shown in Tab. II, EDVD-LLaMA achieves the highest performance across all metrics. Specifically, EDVD-LLaMA obtains a CIDEr score of 2.4579, outperforming VideoLLaMA3 (2.3154), Video-ChatGPT (2.2416), VideoChat (2.1952), and Video-LLaVA (2.1738). Similarly, EDVD-LLaMA achieves the best ROUGE_L (0.7013), BLEU-4 (0.5048), METEOR (0.4236), and CSS (0.7135), consistently surpassing all baselines. These results demonstrate that EDVD-LLaMA generates more accurate, fluent, and semantically consistent explanatory texts for deepfake video detection compared to existing MLLMs.
@@ IV-C Generalization Performance Cross Forgery Methods and Datasets
To further evaluate the generalization ability of different methods across various forgery types, we conducted cross-forgery experiments. As shown in Tab. IV, we performed cross-testing among different forgery types in the ER-FF++set, where models are trained on one type of forgery and tested on the others. The results demonstrate that traditional DVD methods exhibit limited performance when transferring across different types of forgery. In contrast, our proposed method consistently achieves the best average score across forgery methods, indicating superior cross-type generalization capability.
TABLE IV: Cross-forgery evaluation results on the ER-FF++set. Models are trained on one type of forgery and tested on others. Cross Avg. denotes the average score across forgery methods.
Train Methods Test(Intra Dataset——Cross Forgery) Cross Avg.
@@
Face2Face Xception 56.28 - 50.73 49.74 51.43 52.05
@@ IV-D Qualitative results
In Fig. 5, we present a qualitative comparison of EDVD-LLaMA with other baseline models (such as Video-LLaVA [lin2023video], Video-ChatGPT [maaz2024videochatgptdetailedvideounderstanding], VideoChat [li2024videochatchatcentricvideounderstanding], and VideoLLaMA3 [damonlpsg2025videollama3]) on the ER-FF++set. This dataset contains samples of deepfake videos and real videos, where the models are required to determine the authenticity of each video. EDVD-LLaMA demonstrates outstanding performance in detecting deepfake videos through its unique design. The results show that EDVD-LLaMA can accurately identify subtle artifacts in deepfake videos, such as decoupled motion vectors and anomalies in dynamic changes. In contrast, other baseline models often fail to handle these complex scenarios due to insufficient sensitivity to artifacts or incomplete feature extraction, resulting in misjudgments. For example, VideoLLaVA and Video-ChatGPT fail to detect critical artifacts in certain deepfake videos, leading to incorrect predictions, while VideoChat and VideoLLaMA3 occasionally generate false positives when processing real videos. In conclusion, EDVD-LLaMA not only achieves accurate judgments of video authenticity but also provides detailed reasoning, proving the effectiveness and potential of its design. More qualitative results of our model can be found in the supplementary video.